Opened 6 years ago

Last modified 3 years ago

#475 new flight processing

GB12/04, flight day 106/2013, Eaves Wood

Reported by: knpa Owned by: benj
Priority: alpha 4 medium Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/uk/GB12_04-2013_106_Eaves_Wood

Data arrived from ARSF via SATA disk Y on 22/04/2013

GB12/04 Scientific objective:

GB12/04 Priority: a4m

PI: Alan Blackburn

Notes: The IMU failed during this flight. It was re-initialized but there will be a gap in the coverage.

Sensors:

Hawk (Delivered 03/07/2013)
Camera (Delivered 03/07/2013)
Eagle (Delivered 03/07/2013)
LiDAR (Dispatched 16/10/2013)
FW LiDAR (Dispatched 16/10/2013)

Change History (31)

comment:1 Changed 6 years ago by edfi

Starting Navigation processing

comment:2 Changed 6 years ago by tipo

the IMU failure happened between 13:22:9 and 13:32:30, line 3 does not process as it took place between these times.

comment:3 Changed 6 years ago by tipo

Started hyperspectral processing

comment:4 Changed 6 years ago by edfi

Starting LiDAR processing. Checking for navigation issues

comment:5 Changed 6 years ago by tipo

Started RCD processing

comment:6 Changed 6 years ago by tipo

Corrupt pixels found in level 1 files at column 0 (and on band 251 both 0 and 1)

Last edited 6 years ago by tipo (previous) (diff)

comment:7 Changed 6 years ago by tipo

Corrupt pixels found in RCD images at aprox coords
2881, 3772 to 3779, 3776

Last edited 6 years ago by tipo (previous) (diff)

comment:8 Changed 6 years ago by tipo

RCD reprocessed to remove corrupt pixels. Delivery created.

comment:9 Changed 6 years ago by tipo

Waiting for Hyperspectral Data Quality Report

comment:10 Changed 6 years ago by tipo

SCT values:
(Eagle was processed with use_nav = false)

LineEagleHawk
12.59-0.09
21.95-0.04
3N/AN/A
41.70-0.04
50.70-0.04
63.5-0.04
72.4-0.04
82.38-0.04
92.8-0.04
102.0-0.04
112.45-0.02

comment:11 Changed 6 years ago by tipo

Waiting for badpixel generation to be finished.

comment:12 Changed 6 years ago by anhi

RCD Delivery check:

  • Tiff files are not tagged with positional and navigational data

comment:13 Changed 6 years ago by edfi

  • Description modified (diff)

comment:14 Changed 6 years ago by tipo

Tagging fixed.

comment:15 Changed 6 years ago by anhi

RCD Delivery check:

  • Readme exiftool sample output does not have positional or navigational data.

comment:16 Changed 6 years ago by tipo

Regenerated readme file to fix the exiftool example.

comment:17 Changed 6 years ago by anhi

RCD Delivery check:

No problems found.

comment:18 Changed 6 years ago by tipo

Created Hyperspectral delivery.

comment:19 Changed 6 years ago by anhi

Hyperspectral delivery check:

  • viewvector_list.txt files missing.
  • Lines h11 and e11 do not line up with the other lines, maybe can be improved on.

No other problems.

comment:20 Changed 6 years ago by tipo

generated viewvector_list.txt files.
h11 and e11 do not line up well under any SCT values. Noted in readme.

comment:21 Changed 6 years ago by anhi

Hyperspectral changes check:

No problems.

comment:22 Changed 6 years ago by tipo

RCD and Hyperspectral delivered to Alan Blackburn, Lancaster University, 3/7/2013

comment:23 Changed 6 years ago by tipo

  • Description modified (diff)

comment:24 Changed 6 years ago by edfi

Ready to D/C

Due to IMU failure LiDAR data has large errors both vertically and horizontally. Additionally it appears that lines are not consistent within themselves

comment:25 Changed 6 years ago by tipo

Checked LiDAR delivery, no problems found.

comment:26 Changed 6 years ago by tipo

Dispatched LiDAR 16/10/2013

comment:27 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:28 Changed 5 years ago by dac

Started archiving

comment:29 Changed 5 years ago by dac

Finished checking data.
Removed duplicate processed data in processing directory.
Started uploading data to NEODC via rsync.

comment:30 Changed 5 years ago by dac

Data uploaded to NEODC

comment:31 Changed 3 years ago by dac

Looking at LiDAR data for this flight compared to 2012/86, 2012/219b and 2014/83 over the same area there is quite a large horizontal offset (~ 40 m north-west). Due to problems with IMU listed in readme.

Note: See TracTickets for help on using tickets.